OpenStudy (anonymous):

I bet you can't simplify this:

OpenStudy (alexandervonhumboldt2):

simplifying nothing is impossible xD. what is your question?

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" so i agree, you cannot do it

OpenStudy (anonymous):

lol what is the q

OpenStudy (anonymous):

\[Cos/\sin + 1/\sin \times \sin\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

not really clear what that is the equation tool or a screen shot might help

OpenStudy (anonymous):

hello misty

OpenStudy (anonymous):

um, how do I screen shot ^_^?

OpenStudy (alexandervonhumboldt2):

is it \[\frac{ Cos }{ \frac{ \sin+1 }{ \sin*\sin } }\]?

OpenStudy (alexandervonhumboldt2):

to make a screen shot press PrtSc on your keyboard, then open paint, paste your picture and then save file and upload it using Attach file button.

OpenStudy (anonymous):

no it's cos over sin + 1 over sin times sin if you can read that

Nnesha (nnesha):

looks like \[\frac{\color{ReD}{\frac{ \cos }{ \sin +1}} }{ \sin \times \sin }\]

OpenStudy (anonymous):